OpenStudy (anonymous):

what is the the sum of 3 consecutive odd integers is 18 less than five times the middle number. please help!!

OpenStudy (anonymous):

Set up the problem like this: x + x+2 + x+4 = 5(x+2) - 18 where x is the lowest integer, x+2 the middle one, and x+4 the greatest. then simplify: 3x +6 = 5x + 10 -18 3x +6 = 5x - 8 14 = 2x x=7 So the three integers are 7, 9, 11.

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Yw!